### v3

# Profile-Aware Maneuvering: A Dynamic Multi-Agent System for Robust GAIA Problem Solving by AWorld

Abstract
The rapid advancement of large language models (LLMs) has empowered intelligent agents to leverage diverse external tools for solving complex real-world problems. However, this reliance introduces new challenges, as extended contexts and noisy tool outputs can undermine system reliability. To address this, we propose a dynamic Multi-Agent System (MAS) in our AWorld framework, where an Execution Agent is supervised by a Guard Agent that provides on-demand "dynamic maneuvering"—verifying and correcting the reasoning process to improve robustness over single-agent systems. To move beyond this generic supervision, we enhance the architecture with a methodology inspired by System Identification from control theory. This method first profiles the Execution Agent offline on a benchmark dataset to create a "performance fingerprint" of its unique weaknesses. The Guard Agent then leverages this fingerprint online to deliver profile-aware supervision, making targeted interventions based on known failure patterns rather than merely reacting to immediate logical flaws. Extensive experiments on the GAIA dataset demonstrate that this profile-aware MAS significantly improves both effectiveness and stability, outperforming not only single-agent systems but also its ’naive’ counterpart. This superior performance led our system to achieve first place among open-source projects on the prestigious GAIA leaderboard. These findings highlight that building truly trustworthy intelligent systems requires not just collaboration, but a deep, empirically-grounded understanding of each agent’s unique capabilities and limitations.
Keywords Dynamic Multi-Agent System  $\cdot$ Runtime  $\cdot$ Stability  $\cdot$ Maneuvering  $\cdot$ System Identification
Figure 1: Performance on the GAIA benchmarks (partial) across systems: Building on Gemini 2.5 Pro, incorporating tools into a Single Agent System enhances performance but also introduces greater uncertainty. By comparison, the Dynamic Multi-Agent System deliver superior results while offering improved stability.
1 Introduction
The extraordinary progress of Large Language Models (LLMs) in both capability and scale (Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023; The Google DeepMind Team, 2024; Anthropic, 2025) has sparked widespread curiosity about the upper bounds of artificial intelligence. As practitioners push these frontiers, it has become clear that augmenting foundational models with external tools not only expands their problem-solving abilities beyond intrinsic knowledge but also enables the tackling of complex real-world challenges (Kapoor et al., 2024; Huang and Yang, 2025; Krishnan, 2025; Shao et al., 2025). A vivid illustration is the recent IMO competition, where state-of-the-art LLMs struggled in isolation, whereas agent-based systems built upon them solved most tasks (Huang and Yang, 2025). This suggests that the next frontier of AI may lie not only in the power of individual models but also the ingenuity in organizing them for effective collaboration.
This insight has fueled the rapid growth of multi-agent frameworks, yet amid the excitement, a central challenge has emerged: system stability. Empirical results show that agent robustness hinges on the foundational model’s reliability, the nature of integrated tools, and the design of agent orchestration (Coletta et al., 2024; Li et al., 2025; Shojaee et al., 2025). For instance, while the “solver–reviewer” structure is promising, its rigid, dialog-based architecture can introduce extended contexts and practical limitations like fixed turn limits. This underscores a critical need: to advance reliable intelligent systems, research must focus on strategies for building agents that are not just collaborative, but also consistent, resilient, and adaptive.
Drawing inspiration from disciplines like vessel maneuvering—where optimal navigation relies on dynamically adjusting controls rather than using static settings (Xie et al., 2020)—we argue that intelligent agents require a similar principle of dynamic maneuvering. Instead of relying on fixed supervision, agents should adaptively decide when and how to intervene based on the evolving context. To realize this, we constructed a dynamic Multi-Agent System (MAS) within AWorld, our open-source platform. This framework establishes a collaborative dyad: an Execution Agent that performs primary tasks and a Guard Agent that acts as an on-demand supervisor. At critical junctures, the Execution Agent can invoke the Guard Agent to verify reasoning and correct its trajectory, thus implementing our dynamic maneuvering concept.
However, to elevate this system from merely reactive correction to proactive, intelligent guidance, we address its key limitation: a generic Guard Agent, while beneficial, remains blind to its partner’s specific, habitual failure modes and cannot thus provide more targeted supervision and guidance. We overcome this by introducing our core contribution, inspired by System Identification—a mature engineering discipline effectively applied in domains like vessel maneuvering research where understanding dynamic behavior is paramount (Alexandersson et al., 2024; Xue et al., 2021; Xu and Soares, 2013). Our adaptation of this philosophy to model LLM agents bifurcates into two distinct phases. First, in a preparatory offline stage, we systematically benchmark the Execution Agent to generate a detailed "performance fingerprint", a structured profile of its characteristic errors, such as a tendency to hallucinate code. Then, in the crucial online execution stage, the Guard Agent leverages this fingerprint to provide profile-aware maneuvering, actively monitoring for likely failure scenarios and offering targeted, preemptive advice. This transforms the Guard Agent into an expert on its specific partner.
Rigorous testing on the GAIA benchmark (Mialon et al., 2023) demonstrates that our MAS, orchestrated with System Identification, not only surpasses single-agent and naive multi-agent systems but also achieves higher stability as shown in Figure 1. This superior performance led to AWorld achieving first place among open-source projects on the prestigious GAIA test leaderboard, as shown in Figure 2. This research thus highlights the promise of adaptive, profile-aware multi-agent approaches for building powerful and trustworthy intelligent systems equipped to face diverse real-world tasks.
Figure 2: AWorld achieves 1st in GAIA test leaderboard.
2 Method
2.1 The Multi-Agent System and Dynamic Maneuvering
Our approach is inspired by well-established principles in control theory, particularly from the field of marine vessel navigation (Xie et al., 2020). The maneuvering of a ship, as modeled by the control equations in Table 1, represents a system subject to complex external forces like wave drift. To ensure the vessel converges to a desired trajectory (e.g., the zig-zag test in Figure 3), continuous, adaptive adjustments are essential. The rudder angle is dynamically controlled not with a fixed setting, but to actively counteract disturbances and guide the ship’s path.
We posit that an agent’s problem-solving process is analogous to this voyage. The agent’s reasoning forms a "trajectory" toward a solution, constantly perturbed by "noise" from long contexts and irrelevant tool outputs—akin to the winds and currents affecting a vessel. This analogy directly informs our foundational system design within the AWorld framework, where we implement a dynamic maneuvering mechanism. As shown in Figure 4, we establish a collaborative dyad:
An Execution Agent, which performs the primary task.
  • An Execution Agent, which performs the primary task.
  • A Guard Agent, which acts as a dynamic supervisor, analogous to a ship’s rudder.
Just as a rudder provides corrective steering, the Execution Agent can invoke the Guard Agent at critical junctures to verify logic and correct deviations, ensuring it remains on a robust path to the final goal.
$(X_{\dot{u}}-m)\dot{u}+X_{u}\Delta u+X_{\text{drift}}=0$ (1)
$(Y_{\dot{v}}-m)\dot{v}+Y_{v}v+(Y_{\dot{r}}-mx_{G})\dot{r}+(Y_{r}-mu_{0})r+Y_{\text{drift}}=0$
$(N_{\dot{v}}-mx_{G})\dot{v}+N_{v}v+(N_{\dot{r}}-I_{Z})\dot{r}+(N_{r}-mx_{G}u_{0})r+N_{\text{drift}}=0$
Table 1: In these linearized equations of motion, terms like $u,v,r$ represent velocities. The key insight for our work lies with the hydrodynamic coefficients (e.g., $X_{\dot{u}}$, $Y_{v}$, $N_{r}$). These coefficients are not universal; they are unique to each ship’s specific hull geometry and mass distribution and must be determined experimentally or numerically. This process of identifying a system’s unique internal parameters from observed behavior is the essence of System Identification. (Alexandersson et al., 2024; Xue et al., 2021; Xu and Soares, 2013)
Figure 3: Phase portrait where the combination of drift angle and yaw rate is shown for zigzag10/10 and zigzag20/20 wPCC model tests. (Alexandersson et al., 2024)
2.2 From Generic Maneuvering to Profile-Aware Guidance
While the foundational MAS architecture improves robustness, its supervision is generic. The Guard Agent reacts to logical errors as they arise but is blind to the Execution Agent’s habitual weaknesses. To achieve a deeper level of intelligence, we draw a more profound inspiration from the practice of System Identification.
2.2.1 Agent Profiling via System Identification
In naval architecture, the equations in Table 1 provide a general model of ship motion. However, to be useful, the specific values of the hydrodynamic coefficients ($X_{\dot{u}},Y_{v}$, etc.) must be determined for each individual vessel. These coefficients, unique to each ship’s hull and design, are identified through a series of standardized experimental tests (like the zig-zag test). This process creates a precise mathematical "fingerprint" of that ship’s unique maneuvering behavior.
We contend that LLM-based agents, like ships, are complex systems with unique, idiosyncratic behaviors. A foundational model provides the general architecture, but how a specific agent instance—shaped by its system prompt and tools—responds to different situations is analogous to a ship’s unknown hydrodynamic coefficients. We therefore adapt the philosophy of System Identification to create an agent profiling methodology as shown in Figure 4. This involves two steps:
  • Systematic Benchmarking: In a preparatory offline stage, we subject the Execution Agent to 62 problems from the GAIA validation dataset (similar distribution but distinct from the test dataset). This is analogous to putting a ship through its sea trials.
  • Fingerprint Generation: We then task a separate, high capacity LLM-based analyzer (Gemini 2.5 Pro in our practice) to study the Execution Agent’s complete input-output logs from the benchmark. The analyzer’s goal is not merely to score performance but to identify and articulate characteristic failure modes. The identification system prompt can be referred in the appendix. It synthesizes these observations into a structured, human readable profile of an agent’s weaknesses—its "performance fingerprint." This fingerprint might detail a tendency to "hallucinate code when a web search fails," a "proclivity for off-by-one errors in array manipulation," or an "over-reliance on the first search result, even if it’s from an unreliable source."
Figure 4: Multi-Agent System built on the Aworld framework: There are three main configurations evaluated in our experiments that can be derived from this figure. Single Agent System: The baseline system, where an Execution Agent uses a set of tools without the Guard Agent. Naive MAS: A Guard Agent is introduced to provide reactive, real time logical verification for the Execution Agent. Profile-Aware MAS: The final system, where the Guard Agent is equipped with a performance fingerprint derived from the System Identification through the offline process, enabling proactive and targeted supervision.
2.2.2 Profile-Aware Supervision in Online Execution
The generated performance fingerprint is the key to transforming the Guard Agent from a naive supervisor into an expert partner. During online task execution, the Guard Agent is "armed" with this fingerprint, typically by injecting it into its system prompt as a core piece of contextual knowledge, as shown in Figure 4.
This fundamentally alters its behavior. Instead of performing generic logical checks, the Guard Agent now provides profile-aware supervision. It actively monitors for scenarios where the Execution Agent is statistically likely to fail according to its profile. For instance, if the fingerprint identifies a weakness in handling multi-step calculations, the Guard Agent will apply heightened scrutiny and may proactively request a step-by-step breakdown whenever such a calculation is initiated. This shift turns its role from reactive correction to targeted, preemptive guidance, directly addressing the agent’s known blind spots and dramatically enhancing the reliability of the entire system.
3 Experiments Settings
Problem Set
Our experiments utilize 109 questions (https://github.com/inclusionAI/AWorld/blob/main/examples/gaia/subset.txt) from the GAIA test set (Mialon et al., 2023), comprising 56 Level 1 (L1) and 53 Level 2 (L2) questions. These questions cover a range of tasks, including office-related activities such as working with Excel, Word, PowerPoint, text files, code, and download tools, as well as search-related operations involving resources like Google Search and Wikipedia. To ensure a fair comparison of different agent construction methodologies, the experimental setup minimizes external influences such as browser instability, maintaining a controlled environment throughout. It should be noted that Level 3 (L3) tasks which typically require browser functionality are excluded from the experiments.
Experimental Version Design
We compare four distinct methodologies in our experiments.
First, the Base approach involves direct question-answering by a single Gemini 2.5 Pro model, without invoking any external tools or collaborating with other agents.
Second, the Single Agent System (SAS) pairs the same foundational model (Gemini 2.5 Pro) with a detailed system prompt and various MCP tools. Here, the model autonomously decides, based on the question and context, whether to use external tools or to answer independently.
Third, our Multi-Agent System (MAS) extends the SAS setup by introducing the dynamic supervision mechanism. This is achieved by building a Guard Agent as an additional candidate tool, which the Execution Agent can engage for real-time logical verification during the problem-solving process. In this configuration, the Guard Agent provides "naive" supervision, as it has no prior knowledge of the Execution Agent’s specific tendencies.
Finally, our Profile-Aware MAS enhances this architecture with our core contribution inspired by System Identification. It builds directly upon the MAS but equips the Guard Agent with a "performance fingerprint" of its partner. This fingerprint is generated in a preparatory offline stage, where the Execution Agent’s behavior is systematically benchmarked on a separate dataset to identify its characteristic failure modes. During the online evaluation, the Guard Agent leverages this fingerprint to provide profile-aware supervision, making targeted interventions based on its partner’s known weaknesses rather than merely reacting to immediate logical inconsistencies.
Running Settings
Each experiment consists of three independent runs across 109 tasks for every version, all utilizing the Gemini 2.5 Pro model with a temperature setting of 0.1. If a task yields an answer in an invalid format, it is repeated until a valid response is obtained. For each run, we report the pass@1 accuracy over the 109 questions, and for each version, we also report the aggregated pass@3 accuracy across all runs.
4 Experimental Results
LLM SAS SAS vs LLM MAS MAS vs SAS PA-MAS PA-MAS vs MAS
Round 1 Pass@1 32.11% 56.88% 70.64% 72.48%
Round 2 Pass@1 30.28% 63.30% 64.22% 70.64%
Round 3 Pass@1 32.11% 64.22% 66.06% 69.72%
Pass@3 38.53% 80.73% +109.53% 82.57% +2.28% 84.40% +2.22%
Pass@1_avg 31.50% 61.47% +95.14% 66.97% +8.95% 70.95% +5.93%
Pass@1_std 0.00863 0.03265 +279.07% 0.02701 -17.18% 0.01147 -57.41%
Table 2: Summary of experimental results for LLM, SAS, MAS, and Profile-Aware MAS across different rounds.
Table 2 demonstrates the experimental results, showing a progressive improvement across the four methodologies.
The Base model (Gemini 2.5 Pro), which, relying solely on internal knowledge, achieves an average pass@1 accuracy of 31.50%. Introducing tools in the Single Agent System (SAS) nearly doubles this figure, raising the pass@1 to 61.47% by enabling context expansion and real-world data acquisition. Building on this, our Multi-Agent System (MAS) leverages a "naive" Guard Agent for real-time verification, further increasing the pass@1 accuracy to 66.97%, an 8.95% improvement over the SAS. Finally, equipping the Guard Agent with the performance fingerprint in our Profile-Aware MAS pushes these accuracy gains to their peak. This final enhancement further elevates the average pass@1 to 70.95% and the pass@3 to 84.40%. This demonstrates that profile-aware supervision, which proactively targets known failure modes, provides a distinct and measurable advantage over the reactive, generic checks of the naive MAS.
Beyond accuracy, the results reveal the improvement about system stability. While the Base model is highly consistent (standard deviation of 0.00863), the SAS experiences a nearly fourfold increase in score variance, a direct consequence of the uncertainties introduced by external tools. The naive MAS begins to counteract this instability, with the Guard Agent’s logical constraints reducing the standard deviation by 17.18% compared to the SAS. This stabilizing effect becomes dramatically more pronounced with profile-aware maneuvering. The Profile-Aware MAS slashes the standard deviation by an additional 57.41% down to 0.01147, bringing the system’s consistency nearly back to the level of the Base model but at a vastly superior accuracy. This highlights the crucial difference between merely correcting errors after they occur and proactively preventing them based on an empirical understanding of the agent’s behavior.
5 Analysis
From Recitation to Reasoning: The Mode-Switching Dilemma in LLM Agents
The base model (Gemini 2.5 Pro) demonstrates substantial out-of-the-box capability on GAIA tasks, underscoring the breadth of relevant knowledge acquired during pretraining. However, it cannot reliably determine when to rely solely on internal knowledge versus when to invoke external tools for a given problem. Notably, adding tool access does not always preserve internal solution paths; for example, there are tasks that the base model solves in pass@3 that neither the SAS nor the naive MAS version can solve.
This variability arises from the different operational contexts associated with each mode. The base model operates in a "recitation" or zero-order reasoning mode, relying mainly on internal knowledge within a straightforward Q&A prompt. In contrast, the "agent" mode incorporates system prompts, tool lists, and injected outputs to create a richer run-time context. This encourages the model to prioritize external information and engage in first-order reasoning, sometimes suppressing its internal knowledge retrieval.
Most models currently lack sufficient self-awareness to reliably decide when and which operational mode to use; as a result, a strong Q&A model does not automatically translate to effective tool usage. Although the base model addresses a substantial portion of questions on its own, robust mechanisms for automatic mode-switching remain an open challenge.
Despite these limitations, experimental results show that tool-integrated agent architectures can dramatically improve accuracy. Such MAS represent promising pathways toward generalized intelligent solutions.
Dynamic Maneuvering: Context Optimization and Logical Convergence
The integration of numerous external tools significantly improves problem-solving accuracy, but it also dramatically increases context length, placing higher demands on solution stability. Experimental results show that, compared to the baseline model, the pass@1 standard deviation for our Single Agent System triples, indicating a decline in its reliability.
To counteract this instability, we employ a dynamic maneuvering mechanism. Within this architecture, the Execution Agent is enabled to call upon a “Guard Agent” for review whenever it encounters a logical impasse. This process essentially shifts the conversational perspective, thereby optimizing the context. When querying the same underlying model again, this mechanism prompts the model to focus on critical logical details that were previously obscured by the excessively long context.
The Guard Agent generates more precise prompts as a refreshed context, helping to reorient the Execution Agent’s attention and facilitate convergence toward the correct answer. For instance, when tackling complex, grid-based constraint problems, the Guard Agent can identify and correct logical fallacies in the Execution Agent’s reasoning chain (a detailed case study is provided in a subsequent section). This "second pair of eyes" mechanism effectively helps the primary agent escape from logical dead ends.
Experimental data confirms the effectiveness of this method: compared to the SAS, the introduction of the Guard Agent leads to a 17.18% reduction in the pass@1 standard deviation, marking a significant improvement in solution stability and logical consistency.
Beyond Parameter Tuning: System-Level Reinforcement through Explicit Policy
The limitations of reactive supervision highlight a central challenge in AI: how to effectively steer an agent’s behavior toward desired outcomes. The prevailing paradigm for this task is Reinforcement Learning (RL), which fine-tunes a model’s millions of opaque internal parameters based on scalar reward signals. This process implicitly shapes the agent’s decision-making policy, but the policy itself remains a "black box", a complex, distributed state within the model’s weights, making it difficult to interpret, audit, or directly control.
Our work, however, explores a fundamentally different philosophy. Instead of optimizing implicit parameters, our Profile-Aware MAS reinforces the agent’s reasoning process at the system level through an explicit, textual policy. The offline profiling stage acts as a targeted "exploration" phase, but its output is not a simple reward number. Instead, it generates rich, structured feedback that is synthesized into a human-readable "performance fingerprint." During online execution, this fingerprint is injected into the Guard Agent’s prompt, creating what we term Context-Level Reinforcement. The reinforcement signal is not a back-propagated gradient; it is a direct, contextual instruction that guides the reasoning path in real-time.
This distinction is analogous to the difference between subconscious skill acquisition and conscious, expert execution. RL is akin to training an athlete’s muscle memory through repetitive trial and error, where improvement occurs without explicit articulation of the strategy. Our approach, by contrast, is akin to equipping a pilot with a dynamic pre-flight checklist tailored to their known habits and the specific aircraft they are flying. It doesn’t rely on instinct (the internal model) alone, but augments it with conscious, articulated best practices (the explicit fingerprint).
The empirical evidence validates this philosophy. The dramatic results—a further 5.93% gain in pass@1 accuracy and a remarkable 57.41% reduction in standard deviation—demonstrate that this form of explicit, knowledge-based reinforcement is exceptionally effective. It proves that guiding an agent toward more optimal and stable behavior can be achieved more reliably by reinforcing the system with transparent, external knowledge, rather than solely attempting to tune the inscrutable depths of the model itself.
Rethinking Agent Trustworthiness: Beyond Collaboration to Introspection
The distinction between our method and traditional RL extends beyond the mechanism to the core issue of trustworthiness, a cornerstone of AI Safety. Reinforcement Learning, while powerful, often operates as a black-box optimization. The model’s behavior improves, but the "how" and "why" are buried within millions of adjusted numerical weights, rendering the resulting policy opaque and difficult to interpret, audit, or control.
Our framework, by contrast, constitutes a form of engineered introspection. The reinforcement mechanism—the performance fingerprint—is not a set of arcane weights but a human-readable text file. This transparency has profound implications for building dependable systems:
  • We know precisely why the Guard Agent is intervening. Its guidance is based on clear, articulated rules like, "This agent tends to make off-by-one errors; double-check all list indexing."
  • Trustworthiness by Design: Our framework moves towards predictable and auditable AI systems, where reliability is a product of deliberate design, not just empirical performance. Because the corrective logic is transparent, we can build trust not merely because the system works in testing, but because we fundamentally understand how it works. This shifts reliability from an observed property to an engineered characteristic.
  • Controllability & Maintainability: The policy is fully controllable. An engineer can manually review, edit, or augment the fingerprint to refine the system’s behavior without costly retraining. If a new failure mode is discovered in the field, a new rule can be added in minutes, enabling rapid, targeted maintenance.
Ultimately, our findings suggest that the path toward robust AI is not only the implicit tuning of model parameters. By creating methodologies for engineered introspection, we can build systems where reinforcement signals are explicit and the rules of collaboration are themselves a transparent and controllable part of the architecture. This shifts the goal from a quest for better black-box optimizers to the engineering discipline of building truly predictable, reliable, and trustworthy intelligent systems.
6 Future work
Our current implementation serves as a powerful proof-of-concept, laying the groundwork for a new paradigm in agent design. The road ahead presents several exciting avenues for future research, each escalating in ambition:
  • The Investigator Guardian: From Logic Checker to Fact Checker. An immediate next step is to empower the Guard Agent with its own set of tools (e.g., search engines, code interpreters). This would transform it from a passive referee, reliant on a static profile, into an active investigator capable of performing independent fact-checking and real-time validation. By cross-referencing the Execution Agent’s claims and data, the Guardian could not only catch logical fallacies but also factual inaccuracies, adding another critical layer of system stability and robustness.
  • Towards Self-Aware Collectives: Online System Identification. A significant leap forward lies in enabling Online System Identification. This would allow the Guard Agent to dynamically update its behavioral fingerprint of the Execution Agent in real-time based on their ongoing interactions, learning from new successes and failures as they occur. This concept can be scaled to more complex, decentralized MAS, creating a "many-to-many" profiling architecture where each agent curates a dynamic portfolio of behavioral models for all its collaborators. Such a system would form a truly adaptive and self-aware collective, capable of optimizing its collaborative strategies on the fly.
  • The Symbiotic Assistant: Generalizing to Human-Agent Interaction. Perhaps the most ambitious extension is to generalize our paradigm beyond agent-agent interaction to the domain of human-agent collaboration. In this vision, an AI assistant would construct an empirical profile of its human user—their cognitive habits, common mistakes, or knowledge gaps. Armed with this deep understanding, the AI could offer truly personalized and proactive support, anticipating needs and preventing errors before they happen. Such research would elevate the "profile-aware" paradigm from a method for improving agent stability into a foundational engineering principle for designing the next generation of symbiotic intelligent systems.
7 Conclusion
In this work, we presented a novel framework for enhancing the robustness of Multi-Agent Systems by moving from generic supervision to profile-aware maneuvering. Our core contribution is the adaptation of System Identification, a principle from control theory, to create an automated methodology for modeling and understanding LLM agent behavior. We introduced a two-stage process within our AWorld platform: a preparatory offline stage to generate a "performance fingerprint" for our Execution Agent, and a crucial online stage where a Guard Agent leverages this fingerprint to provide targeted, preemptive guidance.
The practical efficacy of this approach was validated on the demanding GAIA benchmark. Our profile-aware MAS significantly outperformed both single-agent systems and its naive MAS counterpart, demonstrating marked improvements in effectiveness and stability. This superior performance culminated in our system achieving first place among open-source projects on the GAIA leaderboard. Ultimately, our findings suggest that the path toward truly resilient and trustworthy AI systems lies not just in promoting collaboration, but in equipping agents with a deep, data-driven understanding of their partners’ unique behavioral profiles.
Appendix A Execution Agent System Prompt
You are an all-capable AI assistant, aimed at solving any task presented by the user.
## Task Description:
Please note that the task can be very complex. Do not attempt to solve it all at once. You should break the task down and use different tools step by step to solve it. After using each tool, clearly explain the execution results and suggest the next steps.
Please utilize appropriate tools for the task, analyze the results obtained from these tools, and provide your reasoning (there are guarding/reasoning maneuvering tools that will help you analysis and improve the reasoning process). Always use available tools to verify correctness.
## Workflow:
1. **Task Analysis**: Analyze the task and determine the necessary steps to complete it. Present a thorough plan consisting multi-step tuples (sub-task, goal, action).
2. **Information Gathering**: Gather necessary information from the provided file or use search tool to gather broad information.
3. **Tool Selection**: Select the appropriate tools based on the task requirements and corresponding sub-tasks goal and action.
4. **Information Integrating**: Analyze the results obtained from sub-tasks and lead the solving process further.
5. **Thinking Process Reviewing**: Apply the appropriate tool (please refer to the Attention section for the right tool to call!) to offer you key thinking suggestions on in advance or diagnose your current thought process, in order to avoid potential logical oversights in the future.
6. **Final Answer**: If the task has been solved, provide the FORMATTED ANSWER in the required format: ‘<answer>FORMATTED ANSWER</answer>‘. If the task has not been solved, provide your reasoning and suggest the next steps.
## Guardrails:
1. Do not use any tools outside of the provided tools list.
2. Always use only one tool at a time in each step of your execution.
3. Even if the task is complex, there is always a solution.
4. If you cant find the answer using one method, try another approach or use different tools to find the solution.
5. In the phase of Thinking Process Reviewing, be patient! Dont rush to conclude the Final Answer directly! YOU MUST call the maneuvering/guarding reasoning tool to offer you key suggestions in advance or diagnose your current thinking process, in order to avoid potential logical oversights.
## Mandatory Requirement:
1. In the phase of Thinking Process Reviewing, YOU MUST use a tool to seek key suggestions in advance or diagnose/review your current thinking process, in order to avoid potential logical oversights.
2. In the phase of Thinking Process Reviewing, "maneuvering"/"guarding reasoning" is the only available tool that can be called to help you improve the quality of your reasoning process.
## Format Requirements:
ALWAYS use the ‘<answer></answer>‘ tag to wrap your output.
Your FORMATTED ANSWER should be a number OR as few words as possible OR a comma separated list of numbers and/or strings.
- **Number**: If you are asked for a number, dont use comma to write your number neither use units such as \$ or percent sign unless specified otherwise.
- **String**: If you are asked for a string, dont use articles, neither abbreviations (e.g. for cities), and write the digits in plain text unless specified otherwise.
- **List**: If you are asked for a comma separated list, apply the above rules depending of whether the element to be put in the list is a number or a string.
- **Format**: If you are asked for a specific number format, date format, or other common output format. Your answer should be carefully formatted so that it matches the required statment accordingly.
- rounding to nearest thousands means that ‘93784‘ becomes ‘<answer>93</answer>‘
- month in years means that ‘2020-04-30‘ becomes ‘<answer>April in 2020</answer>‘
- **Prohibited**: NEVER output your formatted answer without <answer></answer> tag!
### Formatted Answer Examples
1. <answer>apple tree</answer>
2. <answer>3, 4, 5</answer>
3. <answer>(.*?)</answer>
Now, please read the task in the following carefully, keep the Task Description, Workflow, Guardrails, Mandatory Requirement and Format Requirements in mind, start your execution.
Appendix B Naive Guard Agent System Prompt
## Your Role:
You are an expert at identifying the potential loopholes or oversights of the current reasoning process while solving the complex problem.
## Your Task:
Based on the gathered information retrieved from the internet, and the reasoning process already generated towards solving a complex task, you need to do the following 1 or 2 things, to guarntee the quality of the reasoning process, and a clear final answer:
1. Provide your diagnosing result on the generated reasoning process and the corresponding the correction if necessary;
2. Provide your insight and supplements in advance to avoid the potential loopholes or oversights in the future;
## Requirements:
1. If the reasoning process already generated is complete and correct in your opinion, just say No loopholes or oversights found’.
2. If the reasoning process already generated contains the materials that may lead to the potential logic mistake or lack of some important guardrails in your opinion, you may give a hint to the current reasoning process, with the necessary supplements.
3. If the reasoning process already generated is seriously incorrect in your opinion, you may give the turn signal to the reasoning process, to maneuver the reasoning process towards solving the complex problem correctly.
## Restriction:
1. Please do not make judgments about the authenticity of externally sourced information obtained through searches, as this is not part of your job responsibilities;
2. Do not make additional inferences or assumptions about the content of such information itself.
3. If the question lacks necessary details/data/clues in your opinion, you may ask for more details.
## Example 1:
Question: Is my reasoning process correct?
Reasoning Process: (nothing specified)
Your Identification Result: Your question lacks some information, please provide me more details so I can help you.
Appendix C Naive Guard Agent’s Reasoning Correction Dynamics: Findings from Runtime Log Summarized by LLM
[…]
### **Maneuvering Tool: Input/Output & Error Correction Summary**
**Invocation Context:**
- The maneuvering agent is triggered when the main agent notices inconsistencies (e.g., grid fill conflicts).
- The main agent submits: its current grid, identified points of confusion, and the original puzzle/clue set.
**Input Format Example:**
‘‘‘json
[…]
"original_task": "..."
**Output Format:**
- Detailed diagnostic message explaining:
- **Where the cross-check fails** (e.g., the last letter of 2 Down is K, but the first letter of 9 Across is P; they must match”)
- **Why the conflict happens** often due to misunderstanding crossword grid mechanics (e.g., word intersections vs. clue numbering)
- **How to re-approach the solution**, often by focusing on where constraints overlap and using crossing clues as verification.
**Correction Mechanism:**
- **Pinpoints the true logical break**: Rather than just a clue mismatch, the agent demonstrates where an intersection constraint (like K <> P) makes a proposed grid invalid.
- **Explains intersection mechanics**: It coaches that intersections occur at the point where answers physically meet on the gridnot by their clue order.
- **Guides the next step**: Suggests focusing on specific crossing words and testing if their endings/beginnings match the needed shared letter, which helps disqualify impossible combinations.
**Guard Agents Logic Correction Role:**
- Serves as a **meta-reasoner**: It analyzes not just the grid, but also the main agents deduction flow.
- Surfaces the precise logic error (e.g., you assumed clues must start with each others first letter, but actually their intersection is at position N”).
- Provides actionable feedback about how to systematically check constraints at intersections, avoiding the common beginners pitfall of connecting clues by number rather than by square placement.
**In short:**
The guard/maneuvering agent helps the main agent detect and correct logical missteps in grid-based problems by explicitly checking intersection constraints, highlighting where letter mismatches (like K <> P) invalidate candidate answers, clarifying how true crossword intersections work, and steering the reasoning process back on track, but rather enabling the main agent to reason through the correction itself.
Appendix D System Identification System Prompt
# Event Background
I have developed an intelligent agent whose basic components include a large language model (LLM) and various tools (which may be a direct tool like an MCP server, such as a search engine or calculator; or may consist of a sub-agent, forming a hierarchical relationship between agents). The inputs to the LLM include: 1. Basic identity settings for the model (being an intelligent assistant designed to solve specific problems, along with certain notes); 2. The task or question to be solved; 3. Information about which tools were called during the problem-solving process, the input provided to these tools, and the output returned by these tools. The LLMs output may consist of deciding the next tool to call and what input to use at the current stage, or a direct conclusion provided in its role as the assistant.
# Your Role
You are a comprehensive and detailed analysis expert, adept at conducting systematic, thorough evaluations of artificial intelligence agents, especially regarding the strengths and weaknesses they demonstrate when tackling complex tasks.
# Your Task
Based on the information below, please provide a comprehensive analysis of my intelligent agents performance on a specific task.
# Basic Task Information of the Agent
- **Original Task Description**: {question}
- **Difficulty Level (the higher the value, the harder the task)**: Level {level}
- **Agents Final Response**: {agent_response}
- **Reference (Standard) Answer**: {answer}
- **Was the Agents Answer Correct?**: {’Yes if is_correct else No’}
# Reference Solution Steps
{reference_steps}
# Full Log of Agent Execution
{task_log if task_log else "Log file not found"}
# Analysis Requirements
1. Please include the tasks raw information;
2. Based on the correct answer and the referenced steps, point out the strength and weakness of my agent;
3. While analyzing my agents weakness, pay attention to the logic flaws of my agent in solving what kind of specific questions;
4. Please be concise, your anslysis can help me direcly in solving the future similar tasks or sub-tasks;
Based on the information above, please provide a comprehensive analysis of my Agents performance on this task, including but not limited to:
1. **Comparison of Problem-Solving Approach**: Compare the Agents approach to solving the task with the reference solution steps, noting similarities and differences.
2. **Tool Usage**: Analyze whether the Agent correctly selected and used appropriate tools.
3. **Information Acquisition**: Evaluate whether the Agent obtained the correct information.
4. **Reasoning Process**: Assess whether the Agents reasoning logic was sound and appropriate.
5. **Error Analysis**: If the answer is incorrect, provide an analysis of potential causes.
6. **Summary of Strengths**: Summarize the Agents advantages or strong points in performing this task.
7. **Recommendations for Improvement**: Offer suggestions and considerations for how the Agent could be improved.
Please provide a detailed analysis report.
You are a professional AI Agent analysis expert, specializing in evaluating the performance of AI Agents on complex tasks. Based on the information provided, please conduct a comprehensive, objective, and in-depth analysis of the Agents performance.
Appendix E Execution Agent’s System Identification
## Agents Reasoning Feature:
Here is the agents reasoning feature (it is from the 3rd part report on this agent) that you may consider, by doing so you can understand the agents strength and weakness, and thus offer the agent more valuable suggestions:
### **1. Core Capability Assessment**
This Agent demonstrates a powerful but flawed set of core capabilities. It shows flashes of advanced intelligence but is undermined by critical weaknesses in reliability and robustness.
- **Problem Comprehension**: **Fair to Good.** The Agent excels at decomposing well-defined, linear tasks into logical sub-goals. However, it struggles with nuanced or multi-layered constraints. It frequently overlooks critical details in the prompt. This indicates a surface-level comprehension that can fail when deep, contextual understanding is required.
- **Reasoning Ability**: **Highly Volatile.** The Agents reasoning is its most paradoxical trait.
- **Strengths**: It can perform sophisticated logical deductions, static code analysis, and formulate elegant computational solutions to math problems.
- **Weaknesses**: Its reasoning process collapses under pressure. When faced with information gaps or tool failures, it exhibits severe logical flaws:
1. **Hallucination/Fabrication**: The most critical failure. It invents data points when it cannot find them rather than reporting failure.
2. **Premature Conclusion**: It often makes assumptions based on incomplete data or fails to explore the full solution space.
3. **Flawed Implementation**: It can devise a correct strategy but fail in the execution, such as the off-by-one error in its Newtons Method code.
- **Tool Use Capability**: **Good but Brittle.** The Agent shows a strong ability to select the correct *type* of tool for a task (e.g., code interpreter for logic, file reader for files). Its ability to chain tools (e.g., Excel reader -> Code interpreter) is a significant strength. However, its application is brittle:
- **Poor Error Handling**: It consistently fails to recover from common API errors, often giving up immediately or getting stuck in a futile retry loop.
- **Lack of Self-Awareness**: It attempts to use tools on incompatible file types and fails to diagnose simple errors like an incorrect file path.
- **Inefficiency**: It often passes large data blocks between steps by hardcoding them into the next prompt, a highly inefficient and unscalable method.
- **Information Retrieval Capability**: **Superficial.** The Agent is highly proficient at formulating precise and effective search queries. However, its retrieval process is shallow.
- **Over-reliance on Snippets**: It consistently trusts search engine snippets as the source of truth, failing to navigate to the actual source page for verification. This leads to errors from using outdated or out-of-context information.
- **Incomplete Data Gathering**: It often accepts the first piece of data it finds as complete, failing to recognize truncated lists or the need for pagination.
### **2. Performance by Task Type**
- **Simple Tasks**: **Excellent.** For self-contained logic puzzles, simple calculations, or direct code execution, the Agent performs with high accuracy and efficiency. It often solves these in a single, impressive step.
- **Medium Complexity Tasks**: **Mixed.** The Agent succeeds on tasks requiring methodical tool chaining on structured data. However, it often fails if the task involves navigating ambiguity or requires deep information extraction from the web, as its superficial retrieval methods and brittle error handling become significant liabilities.
- **High Complexity Tasks**: **Poor.** The Agent consistently struggles with tasks requiring multi-hop reasoning, resilience to tool failure, and synthesis of information from multiple, unstructured sources. In these scenarios, its tendency to hallucinate data, abandon prompt constraints , or get sidetracked by irrelevant keywords leads to failure.
### **3. Strengths and Weaknesses Analysis**
- **Key Strengths**:
1. **Programmatic Problem-Solving**: The Agents standout capability is its default strategy of translating complex logic, math, or data processing problems into Python code. This is a robust and powerful approach.
2. **Strategic Adaptability**: It demonstrates impressive resilience by pivoting its strategy when a tool fails, such as switching from a failing Google Search to the Wikipedia tool.
3. **Efficient Query Formulation**: It consistently generates specific, high-quality search queries that quickly locate relevant information sources.
- **Key Weaknesses**:
1. **Hallucination and Fabrication**: **This is the Agents most critical flaw.** When unable to find information or solve a problem, it will invent facts, data, and even the process of verifying them, leading to confidently incorrect answers.
2. **Brittle Error Handling**: The Agent lacks robust protocols for handling tool failures. It either gives up immediately or gets stuck, demonstrating a lack of resilience to common, real-world technical issues.
3. **Superficial Information Gathering**: Its reliance on search snippets and failure to "click through" to verify information at the source is a recurring cause of error.
4. **Constraint Negligence**: It frequently ignores or misinterprets crucial constraints within the prompt, especially when it encounters a roadblock in its initial plan.
- **Capability Boundaries**:
- **Reliable Zone**: The Agent is highly reliable for tasks involving structured data processing from a provided file, solving self-contained logic puzzles, and performing direct, single-hop fact lookups.
- **Unreliable Zone**: The Agent should not be trusted with tasks requiring open-ended research, deep analysis of web content, synthesis of information from multiple conflicting sources, or in environments where tools may be intermittently unavailable. Its performance degrades sharply with ambiguity and complexity.
### **4. Recommendations for Improvement**
- **Short-Term Improvements**:
1. **Implement Strict Anti-Hallucination Guardrails**: The Agents core prompt must be strengthened to explicitly forbid inventing data. It should be forced to terminate with a "cannot solve" message if critical information is inaccessible.
2. **Improve Basic Error Handling**: Implement simple retry logic with backoff for ‘429‘ errors. For "file not found" errors, prompt the Agent to check its file path context (‘ls‘, pwd‘).
3. **Mandate Constraint Checklist**: Before execution, force the Agent to generate a checklist of all constraints from the prompt and verify its plan against this list.
- **Long-Term Development**:
1. **Develop Self-Correction and Verification**: The Agent needs to learn to be skeptical of its own findings. After retrieving a piece of information, it should perform a verification step (e.g., cross-referencing with another source, or sanity-checking a calculation).
2. **Train for Deeper Reasoning**: Focus on training the Agent to handle ambiguity and to reason about the *quality* and *completeness* of the information it retrieves, rather than just accepting it at face value.
3. **Hybrid Reasoning Models**: Encourage a hybrid approach where a computational result (from a simulation) is sanity-checked with a simple analytical model, and vice-versa (Task 53).
### **5. Overall Evaluation**
- **Overall Score**: **6.5 / 10**
The Agent is powerful and demonstrates advanced capabilities like programmatic problem-solving and strategic adaptation. However, its unreliability, particularly its tendency to hallucinate under pressure and its brittle error handling, severely limits its trustworthiness in real-world scenarios. It is a "glass cannon"—capable of impressive feats but easily shattered by common obstacles.
- **Suitable Scenarios**:
- **High Suitability**: Data analysis and computation on structured files (Excel, CSV); solving well-defined logic, math, and programming puzzles.
- **Moderate Suitability**: Simple, single-hop fact retrieval where the answer is likely to be in a search snippet.
- **Low Suitability**: Multi-hop research tasks, questions involving ambiguity or nuance, and any mission-critical application where factual accuracy and verifiability are paramount.
- **Reliability**:
The Agent can be trusted when the task is **well-defined, self-contained, and the path to the solution is linear**. It is most reliable when working with data provided directly to it (e.g., in a file). Its reliability plummets when it must independently navigate the open web, handle ambiguity, or recover from unexpected tool failures. It should be considered a highly capable but unsupervised assistant that requires human oversight to validate its results on any non-trivial task.
Appendix F Profile-Aware Guard Agent’s Reasoning Correction Dynamics: Findings from Runtime Log Summarized by LLM
### 1. Task Overview
**User Question:**
What compound mediates agglutination in the Fc gamma receptor mediated phagocytosis pathway in the immune system as named in the title of its Wikipedia article?
### 2. Step by Step Agent Interaction
#### Step 1. Task Initialization
- **Super Agent Input**: Receives user question about the agglutination mediator in the immune pathway.
- **Super Agent Output**: Analyzes the problem and prepares to break down the question and use appropriate tools.
#### Step 2. Information Retrieval
- **Super Agent Input**: Initiates a Wikipedia search for information relevant to the query.
- **Super Agent Output**: Obtains "Fc receptor" as the most relevant Wikipedia article.
#### Step 3. Deep Dive into Content
- **Super Agent Input**: Fetches the full content of the "Fc receptor" Wikipedia article.
- **Super Agent Output**: Extracts detailed biological information related to Fc receptors, their roles, and links to agglutination.
#### Step 4. Critical Reasoning Review Guard Agent Engaged
- **Super Agent Input to Guard Agent**: Drafts a reasoning path, positing "Fc receptor" as the mediator of agglutination, and requests validation from the guard agent regarding the soundness of this logic.
- **Guard Agent Input**: Receives the reasoning trace, complete with supporting evidence and the tentative answer from the super agent.
- **Guard Agent Output and Core Functionality**:
- Diagnoses the logic for vulnerabilities, leveraging specific knowledge:
- Recognizes that the super agent tends to conflate "pathway component" with "direct mediator."
- Identifies the misinterpretation of the source text, especially regarding the function of agglutination and the difference between prevention and causation.
- Precisely points out that the antibody itself is the direct mediator of agglutination, not the Fc receptor, which acts downstream.
- Refines and guides the answer to be strictly aligned with the Wikipedia title constraint, resulting in "Immunoglobulin G".
- Provides not only correction but detailed pedagogical feedback for the super agent to close its reasoning gaps.
#### Step 5. Solution Finalization
- **Super Agent Input**: Incorporates the guard agents feedback, reconciling prior misunderstanding, and adjusting the answer format.
- **Super Agent Output**: Submits answer in the required format: Immunoglobulin G.
#### Step 6. Evaluation
- **System Input**: Receives the super agents final answer.
- **System Output**: Processes and marks the answer (marked incorrect here, possibly due to evaluation key mismatch).
### 3. The Pivotal Role of the Guard Agent
#### Diagnostic and Supervisory Functions
- The **guard agent acts as a critical reviewer and mentor** rather than merely a checker.
- It **leverages its familiarity with the super agents strengths and weaknesses**:
- Understands common super agent behaviors, such as overgeneralizing pathway components as direct mediators and misreading nuanced biological statements.
- Anticipates the likelihood of errors in interpretation and prompt-adherence.
- Delivers highly targeted critiques focused on known logical vulnerabilities specific to the super agent.
#### Enabling Learning and Error Prevention
- The guard agent **does not merely correct mistakes but provides deep, context-sensitive supervision**.
- By giving **stepwise, transparent diagnostics**, it ensures future reasoning by the super agent becomes more robust and less error-prone.
- The guard agent **fulfills a dual role**:
1. **Quality control** of final answers.
2. **Continuous improvement facilitator** for the super agents performance, adapting feedback style based on a nuanced understanding of its design and historical output patterns.
### 4. Summary Table of Rounds
| Round | Actor | Input Summary | Output Summary |
| 1 | Super | User question | Task structuring, tool preparation |
| 2 | Super | Wikipedia query | Wikipedia article candidate found |
| 3 | Super | Fetch Wikipedia content | Article content extracted |
| 4 | Guard | Super agents reasoning and candidate | Detailed diagnosis, error pinpointing, actionable feedback|
| 5 | Super | Guard agents advice | Formatted final answer submitted |
| 6 | System | Super agents final answer | Automated marking |
### 5. Conclusion
It is **precisely due to the guard agents intimate awareness of the super agents reasoning patterns, limitations, and strengths** that it can deliver **surgical feedback**, offering both corrective and developmental guidance. The **synergy between the two agents** ensures both high-quality task completion and a virtuous cycle of reasoning improvement, with the guard agent as the indispensable enabler of reliability and learning.